NASA Logo, National Aeronautics and Space Administration

National Aeronautics and Space Administration

Goddard Space Flight Center

Astrophysics Science Division | Sciences and Exploration

This website is kept for archival purposes only and is no longer updated.

Skip Navigation
Mission Critical Multi-Media Gallery Hubble News Mission Updates Launch Info

  Mission Updates :: Mar 6 - Flight Day 6 (EVA 3)

Mission Updates
Mission Chronicles
Where Is Hubble Now?
 Related material for Flight Day - REPLACE POWER UNIT
  Click here for Still Images.
  Click here for Video & Audio Clips.
  Click here for Mission Chronicles - the lighter side.

Goddard Space Flight Center
The Hubble Space Telescope Project
Hubble Status Report

The SpacewalkerWed | Mar. 6, 2002 - 1:45 pm EST
The HST Mission Operations Manager has reported that the Space Telescope Operations Control Center (STOCC) has completed over 14 hours of power-on, reconfiguration, and checkout of the Hubble Space Telescope following its complete power-down earlier today.

All HST temperatures have returned to nominal levels. The new Solar Arrays, Diode Box Assemblies, and Power Control Unit have all passed their aliveness and functional tests. The Pointing Control System hardware was successfully powered on and is performing within specification, including the newly-installed Reaction Wheel Assembly #1.

The new Solar Arrays have been successfully slewed to demonstrate they are capable of moving to optimal sun angles during normal operations.

The Science Instruments and Data Handling Computer were successfully recovered from their respective "safe modes". The Faint Object Camera (FOC) remains in "safe mode" for warm-up prior to replacement with the Advanced Camera for Surveys (ACS). All-in-all it has been a tremendously busy and successfully day, following an amazing effort by all during EVA3.

Goddard Space Flight Center
The Hubble Space Telescope Project
Hubble Status Report

Old PCU being removedWed | Mar. 6, 2002 - 1:45 pm EST
The new Power Control Unit (PCU) installed during EVA3 has now passed its lengthy Functional Test with flying colors.

The new PCU, in combination with the new Solar Arrays, provide a major upgrade to the reliability and output of the HST power system and ensures it will support the needs of the full-complement of science instruments, as they are upgraded in this and the next Servicing Mission.

The HST gyros have now also been powered back on successfully (except for gyro 5 which remains failed, as expected). The current Science Instruments are now all receiving power in their standard "safe modes" and will shortly be brought up to various low-level operate or hold modes. All continues to go extremely well.

Johnson Space Center
Space News :: Latest Items

In the bayWed | Mar. 6, 2002 - 10:42 am EST
Spacewalkers John Grunsfeld and Rick Linnehan are back on board Columbia after concluding a 6 hour 48 minute EVA to install a new power unit in the Hubble Space Telescope.

This third spacewalk of the STS-109 mission began at 2:28 a.m. central time and concluded at 9:16 a.m.

A fourth spacewalk will be conducted tomorrow by Jim Newman and Mike Massimino to install a new camera in the orbiting telescope, the Advanced Camera for Surveys.

The crew will begin an abbreviated 7 1/2 hour sleep period just before 2:30 p.m. central time.

Goddard Space Flight Center
The Hubble Space Telescope Project
Hubble Status Report

Hubble's power returnsWed | Mar. 6, 2002 - 9:57 am EST
Following its successful installation on HST by astronauts John Grunsfeld and Richard Linnehan, the replacement Power Control Unit (PCU) passed its Aliveness Test. The Functional Test is about to begin.

Johnson Space Center
Space News :: Latest Items

Old PCU being removedWed | Mar. 6, 2002 - 9:45 am EST
About 5 1/2 hours into today's spacewalk to install a new power control unit, the Hubble Space Telescope's heartbeat was restored at 8:02 a.m. central time.

That heartbeat indicated that telemetry from the telescope was being received on the ground, allowing commands to be sent to the Hubble.

At 8:41 a.m. ground controllers at the Space Telescope Operations Control Center confirmed that aliveness tests of the new power control unit were successful. They will now begin the lengthy process of restoring power to each of the Hubble's various components.

Today's spacewalk, which began at 2:28 a.m., is being conducted by John Grunsfeld and Rick Linnehan. Grunsfeld and Linnehan also conducted the first of five scheduled spacewalks for this mission.

Tomorrow, astronauts Mike Massimino and Jim Newman will conduct the fourth spacewalk of this mission to install a new camera, the Advanced Camera for Surveys, in the orbiting observatory.

Wed | Mar. 6, 2002 - 9:02 am EST
PCU has been changed-out. Telemetry has been restored and once again Hubble has a "heartbeat". Aliveness tests have been so far successful.

Wed | Mar. 6, 2002 - 4:55 am EST
PCU powerdown complete. Astronauts Grunsfeld and Linnehan will begin removing the PCU'S connectors, so that it can be removed. A new one will take its place.

Wed | Mar. 6, 2002 - 3:39 am EST
Wednesday's spacewalk by John Grunsfeld and Richard Linnehan began at 2:28 a.m. CST, after a delay caused by a water leak in Grunsfeld's suit. The astronauts will replace the Hubble Space Telescope's Power Control Unit during the spacewalk, the third of five during the STS-109 mission to upgrade the orbiting observatory. CLICK HERE FOR MORE IMAGES...

Beginning today's spacewalkWed | Mar. 6, 2002 - 2:04 am EST
A water leak in John Grunsfeld's spacesuit has delayed the start of this morning's spacewalk. Grunsfeld will use another spacesuit, which is being resized to fit him. The spacewalk is expected to begin about 2:30 a.m. CST. Grunsfeld and fellow spacewalker Richard Linnehan will change out the Power Control Unit of the Hubble Space Telescope.

Controllers at the Space Telescope Operations Control Center in Green Belt, Md., will power down Hubble for the first time ever to enable Grunsfeld and Linnehan to replace the PCU. Hubble will be powered back up after the new PCU is installed.

 More material about this mission day
JSC Status Report #11
JSC Status Report #12
Commander Grunsfeld's Notes from Space #4
Latest ::
Post-Mission ::
Late Apr
Mid Apr
Early Apr
Late Mar
Mission ::
Mar 12
Mar 11
Mar 10
Mar 9
EVA 5 : Mar 8
EVA 4 : Mar 7
EVA 3 : Mar 6
EVA 2 : Mar 5
EVA 1 : Mar 4
Mar 3
Mar 2
Mar 1
Pre-launch ::
Feb Week 4
Feb Week 3
Feb Week 2
Feb Week 1
Preparation ::
Late Jan
Mid Jan
Early Jan
Late Dec
Mid Dec
Early Dec
Late Nov
Mid Nov
Early Nov
Late Oct
Mid Oct
Early Oct
Glossary | FAQ | Links
| Page Last Updated: March 28, 2012 |
Links FAQ Glossary Home